OpenStudy (kellyspeakslouder):

f(x) = 4x + 7, g(x) = 3x2 Find (fg)(x). A. 12x2 + 21x B. 3x2 + 4x + 7 C. 12x3 + 21x2 D. 12x + 21

OpenStudy (kellyspeakslouder):

@jim_thompson5910

OpenStudy (danjs):

That means to multiply the two functions... \[\large (f*g)(x) = f(x) * g(x) = (4x+7)*(3x^2)\]

OpenStudy (kellyspeakslouder):

what I got isn't one of the options .-.

OpenStudy (kellyspeakslouder):

@DanJS

OpenStudy (danjs):

what did you get?

OpenStudy (kellyspeakslouder):

21x squared + 4x

OpenStudy (kellyspeakslouder):

@DanJS

OpenStudy (kellyspeakslouder):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):